Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-11-03

Observers: Georgios Dimitriadis, Miika Pursiainen
Support Team: Southampton
TAT on duty: Regis Cartier, Joe Lyman

All following times in UT

The night IS photometric

Weather conditions:
Start - No clouds, very low (5 m/s) wind, seeing below 1”
Middle - No clouds, low (5-10 m/s) wind, seeing below 1”
End - Clear skies, no wind, seeing below 1”

Calibrations:

PESSTO calib
PESSTO bias
PESSTO_SkyFF_eve_gr, last z exposure skipped, too dark
PESSTO_SkyFF_mor_UVi

Science:
23:45 We start with EFOSC2
23:50 Image analysis and telescope focus
00:15 AT2016hnq, classification, seeing ~0.8”,low S/N, seems like a galaxy spectrum, ground team may do something better
00:30 AT2016hqf, classification, seeing ~0.9”, SN Ia (99aa-like), at peak, z=0.055
00:45 Feige 110, spectroscopic standard star
01:00 PG2336, photometric standard field, UBVRgriz, seeing ~0.9”
01:10 LSQ15adm, follow up, Vgri, seeing ~0.85”
02:10 Re-focusing the telescope + image analysis
02:30 There seems to be some elongation at the images, problem is fixed with focusing and image analysis, but TO suggested to focus the telescope before each science exposure and we agreed.
02:40 SN2015L, follow up, Bgri, seeing ~0.9”
03:20 SN2009ip, follow up, BR, seeing ~0.8-0.9”
04:05 PG2336, photometric standard field, UBVRgriz, seeing ~0.8”
04:15 PS15cwo, follow up, gri, seeing ~0.8-1” (we split the OB to the three filters, with refocusing before each filter exposure)
06:15 ASASSN-15og, follow up, BVgri, seeing ~0.6-0.9”
07:10 Internet down at the NOB
07:30 RU152,photometric standard field, UBVRgriz, seeing ~0.9”
07:40 AT2016htn, classification, seeing ~0.9”, SN Ia, -5d, z=0.054
08:00 AT2016htm, classification, seeing ~0.85”, SN Ia, SNID gives equally good fits from -6 to peak, at z=0.043 (due to internet problems at the NOB, we cannot see the ATLAS LC, for a better estimate of the phase we leave it to the ground team) 
08:15 L745a, spectroscopic standard star, seeing ~0.9”
08:45 AT2016hrc, classification, seeing ~0.9”, wrong slit position, star spectrum (due to lack of internet, we didn’t have the best finder chart…)
09:00 AT2016hnl, classification, seeing ~0.8”, SN Ia, +10d, z=0.032
09:15 End of night, still Internet down, will update in the morning

NOTE: Night time TO suggested that the day time TO should run some tests in the morning, regarding the possible persistent defocus of the telescope, we will update on the outcome on tomorrow’s night report
Comments